HETAV PATEL

Grade 11,

Two parallel tall buildings are x distance apart.A boy throws a ball horizontally with a velocity 10ms-1 which falls 10m below the opposite building.Then x is,

Two parallel tall buildings are x distance apart.A boy throws a ball horizontally with a velocity 10ms-1 which falls 10m below the opposite building.Then x is,

Grade:12

2 Answers

Arun
4 years ago
Height = h= 55-10.9 
= 44.1
Distance= s= 180 m
time taken = t
so, h=1/2 gt²
or t= √2h/g
√2 x 44.1/9.8
=√88.2/9.8
=√9 = 3
 
S= velocity x t
180= velocity x 3
velocity =180/3 = 60 m/s
 
Hope it helps
Vikas TU
4 years ago
 
Dear student 
There is no relation to question , in above ans.
Please give the data for time or acceleration , I think more data is required, 
Please attach an image if possible, 
Good Luck